# Diff-RNTraj: A Structure-aware Diffusion Model for Road Network-constrained Trajectory Generation

### v1

@@ V-B Evaluated Metrics
Besides evaluating the overall trajectory distribution, we design a novel metric, called Road Segment Connectivity (RSC), to assess the spatial validity of the generated trajectories. RSC measures the reachable proportion of adjacent points within generated trajectories. We determine the road segment connectivity by UTGraph $\mathcal{G}_{\tau}$ since it is reflected by the real vehicle trajectory. For a generated RNTraj $\hat{\tau}_{RNTraj}=\langle\hat{q}_{1},\cdots,\hat{q}_{T}\rangle$, its RSC is calculated as follows:
@@ V-C Baselines
JSD-TD ($\downarrow$) JSD-SD ($\downarrow$) JSD-GPD ($\downarrow$) JSD-RS ($\downarrow$) RSC ($\uparrow$) JSD-TD ($\downarrow$) JSD-SD ($\downarrow$) JSD-GPD ($\downarrow$) JSD-RS ($\downarrow$) RSC ($\uparrow$)
@@
Diff-RNTraj 0.0221 0.0007 0.0281 0.0456 91.01% 0.1165 0.0009 0.0856 0.0399 92.80%
V-D Setting
We implement Diff-RNTraj by PyTorch framework with the Adam optimizer for 30 epochs and a batch size of 256. The learning rate is set to 1e-3 and is halved every 3 epochs. For the denoising network, we set the channel dimensional $C=512$, the positional encoding dimensional $F=512$, the layer of each block $m=4$ and the kernel size to 3 in dilation convolution with padding followed by [25]. We adopt the quadratic schedule for the variance schedule: $\beta_{n}=(\frac{N-n}{N-1}\sqrt{\beta_{1}}+\frac{n-1}{N-1}\sqrt{\beta_{N}})^{2}$ with the minimum noise level $\beta_{1}=0.0001$ and the maximum noise level $\beta_{N}=0.02$. We search the number of the diffusion steps $N$, residual dilation convolution layers $L$ and the embedding dimension $d$ of pre-trained road segment representation within the parameter space ($N\in\{50,100,200,500,1000\}$, $L\in\{4,8,10,15,20\}$, $D\in\{16,32,64,128\}$).
@@ V-E Experiments Results
Firstly, rule-based methods can effectively maintain the spatial validity of trajectories. However, the generated trajectory data struggle to capture real movement patterns, resulting in significant disparities compared to the characteristics of the original trajectories. This indicates that describing complex human travel patterns using simple rules is challenging.
@@ V-G Hyperparameter Analysis
V-H Road Segment Simularity
To assess the confidence of our model in generating discrete road segments, we conduct an experiment that generates 2,000 trajectories with a length of 20 by $N$ steps reverse denoising process, i.e., we obtain the road segment representations $\hat{X}_{e}$ for the 40,000 trajectory points. For each generated road segment representation, we calculate its cosine similarity with the pre-trained road segment representations and record the top 10 most similar road segments. Then, we calculate the average similarity across all trajectory points for different top-$k$ similarity values ($k=1,\ldots,10$).
@@ V-I Visualization Analysis
(a) Heatmap of real trajectory data
(b) Heatmap of generated data based on our model
(c) Heatmap of generated data based on Diff-Bit
(d) Heatmap of generated data based on EETG+HMM
Figure 9: Heatmap of generated trajectory on Porto dataset.
(a) Heatmap of real trajectory data
(b) Heatmap of generated data based on our model
(c) Heatmap of generated data based on Diff-Bit
(d) Heatmap of generated data based on EETG+HMM
Figure 10: Heatmap of generated trajectory on Chengdu dataset.
@@
V-I1 Trajectory Point Distribution Visualization
Figure 11: Single trajectory with lengths of 20, 30, 40, and 50 visualizations on the Chengdu dataset.
V-I2 Single Trajectory Visualization
We randomly generate trajectories of different lengths (20, 30, 40, and 50) and plot the GPS points on the road network in Figure 11. We observe that these trajectories exhibit patterns consistent with common driving behaviors and show a preference for major roads. Additionally, the vehicles consistently move forward along the roads without any instances of reversing, thus validating that using RNTraj to represent trajectory is suitable.
V-J Scability Analysis
@@
Figure 12: The performance of trajectory prediction on different datasets by mixing real and generated trajectories.
V-K Downstream Task
Due to privacy concerns, the limited available trajectory data often hinders the progress of downstream tasks since they require large-scale data. In this experiment, we take trajectory prediction as the downstream task to evaluate the quality of the generated trajectory. Specifically, we aim to predict the next road segment, given the historical road segment sequence.
We divide the real trajectory dataset into training, validation, and testing sets with a ratio of 6:2:2. In the training set, we mix real and generated trajectory data, varying the proportion of real data of 0%, 25%, 50%, 75%, and 100%. We train the trajectory prediction model based on LSTM. The trajectory prediction results of different generated models are shown in Figure 12. We observe that the trajectory prediction performance of Diff-RNTraj closely approximates real data. It is worth noting that even when entirely relying on generated results (i.e., with 0% real data), our model still achieves satisfactory performance. This demonstrates the effectiveness of our model for trajectory-based applications.
VI Conclusion

### v2

@@ 5.2 Evaluated Metrics
Besides evaluating the overall trajectory distribution, we design a novel metric, called Road Segment Connectivity (RSC), to assess the spatial validity of the generated trajectories. RSC measures the reachable proportion of adjacent points within generated trajectories. We determine the road segment connectivity by UTGraph $\mathcal{G}_{\tau}$ since it is reflected by the real vehicle trajectory. For a generated RNTraj $\hat{\tau}_{RN}=\langle\hat{q}_{1},\cdots,\hat{q}_{T}\rangle$, its RSC is calculated as follows:
@@ 5.3 Baselines
  • Markov [60]. Markov treats the road segments as states and constructs a transition matrix to capture the first-order transition probabilities between these road segments.
@@
JSD-TD JSD-SD JSD-GPD JSD-RS RSC JSD-TD JSD-SD JSD-GPD JSD-RS RSC
@@
[…]
5.4 Setting
We implement Diff-RNTraj [^4] https://github.com/wtl52656/Diff-RNTraj by PyTorch framework with the Adam optimizer for 30 epochs and a batch size of 256. The learning rate is set to 1e-3 and is halved every 3 epochs. For the denoising network, we set the channel dimensional $C=512$, the positional encoding dimensional $F=512$, the layer of each block $m=4$ and the kernel size to 3 in dilation convolution with padding followed by [35]. We adopt the quadratic schedule for the variance schedule: $\beta_{n}=(\frac{N-n}{N-1}\sqrt{\beta_{1}}+\frac{n-1}{N-1}\sqrt{\beta_{N}})^{2}$ with the minimum noise level $\beta_{1}=0.0001$ and the maximum noise level $\beta_{N}=0.02$. We search the number of the diffusion steps $N$ and residual dilation convolution layers $L$ within the parameter space ($N\in\{50,100,200,500,1000\}$, $L\in\{4,8,10,15,20\}$). To obtain the pre-trained road segment representation E using the Node2vec algorithm, we generate 100 random paths for each node, each with a length of 80. We set the contextual window size of 10, perform a total of 1000 iterations, and search the number of embedding dimensions $D$ across the parameter space $\{16,32,64,128\}$.
@@ 5.5 Experiments Results
Firstly, rule-based methods RWRN and Markov can effectively maintain the spatial validity of trajectories. However, the generated trajectory data struggle to capture real movement patterns, resulting in significant disparities compared to the characteristics of the original trajectories. This indicates that describing complex human travel patterns using simple rules is challenging.
@@ 5.7 Hyperparameter Analysis
5.8 Effectiveness Analysis of the Generated Road Segment Representation
@@
To assess the effectiveness of our model in generating road segments, we conduct an experiment to check whether our generated road segment representations can be aligned with pre-trained representations. Specifically, we generate 2,000 trajectories with a length of 20 by $N$ steps reverse denoising process, i.e., we obtain the road segment representations $\hat{X}_{e}$ for the 40,000 trajectory points. For each generated road segment representation, we calculate its cosine similarity with the pre-trained road segment representations and record the top 10 most similar road segments. Then, we calculate the average similarity across all trajectory points for different top-$k$ similarity values ($k=1,\cdots,10$).
@@ 5.9 Visualization Analysis
(a) The real trajectories.
(b) The trajectories generated by our model.
(c) The trajectories generated by Diff-Bit.
(d) The trajectories generated by EETG+HMM.
Figure 9: Heatmaps of trajectories in Porto.
(a) The real trajectories.
(b) The trajectories generated by our model.
(c) The trajectories generated by Diff-Bit.
(d) The trajectories generated by EETG+HMM.
Figure 10: Heatmaps of trajectories in Chengdu.
5.9.1 Trajectory Point Distribution Visualization
Figure 11: Visualizations of generated trajectories with lengths 20, 30, 40, and 50 in Chengdu.
5.9.2 Trajectory Visualization
We randomly generate trajectories of different lengths ($T$ = 20, 30, 40, and 50) and plot the GPS points on the road network in Figure 11. We observe that these trajectories exhibit patterns consistent with common driving behaviors and show a preference for major roads. Additionally, the vehicles consistently move forward along the roads without any instances of reversing, thus validating that using RNTraj to represent trajectory is suitable.
5.9.3 Trajectory Diversity Visualization
To explore the diversity of the generated trajectories, we randomly select three origin-destination pairs from the Chengdu dataset and plot the generated. As shown in Figure 12, our model generated multiple paths for each origin-destination pair, ensuring a level of diversity and avoiding overly deterministic movement patterns. The majority of these trajectories followed the major roads, with transitions to alternative routes as they approached the destination, reflecting common driving behavior.
However, it is challenging to determine whether certain generated trajectories are abnormal, such as whether the brown trajectory in the middle of Figure 12 represents a detour. This limitation is primarily due to the model’s emphasis on modeling the overall distribution and its lack of consideration for conditional information guidance.
Figure 12: Visualization of trajectory diversity between the same origin and destination regions in Chengdu.
@@ 5.9.3 Trajectory Diversity Visualization
5.10 Scability Analysis
@@
Figure 13: The performance of trajectory prediction on the Porto dataset by mixing real and generated trajectories.
Figure 14: The performance of trajectory prediction on the Chengdu dataset by mixing real and generated trajectories.
5.11 Downstream Task
In this experiment, we take trajectory prediction as the downstream task to evaluate the quality of the generated trajectory. Specifically, we aim to predict the next road segment, given the historical road segment sequence. We divide the real trajectory dataset into training, validation, and testing sets with a ratio of 6:2:2. In the training set, we keep the overall number of trajectories constant, gradually decrease the ratio of the real trajectories, increase the ratio of the generated trajectories, and train the trajectory prediction model. As the ratio of real trajectories decreases, if the prediction performance change is smaller at the test set, it proves that the spatial transfer patterns of the generated trajectories are closer to the real trajectories, and the generated trajectories are more realistic.
Specifically, we choose LSTM as the trajectory prediction model and vary the proportion of real data in the training set to 0%, 1%, 2%, 5%, 10%, 20%, 50%, and 100%. The trajectory prediction results of different generated models are shown in Figure 13 and 14. We observe that Diff-RNTraj demonstrates superior trajectory prediction performance compared to EETG and Diff-Bit, proving that the real trajectories can be replaced by our generated trajectories. It is worth noting that when the proportion of real data is less than 10%, or even completely relying on generated results (i.e., with 0% real data), our model still achieves satisfactory performance, with an average improvement of 3.18% and 3.81% on the Porto and Chengdu datasets compared to Diff-Bit. This is because Diff-RNTraj incorporates the pre-trained road segment representation module to model the user’s road transfer pattern, enabling the generation of realistic trajectories. This demonstrates the effectiveness of our model for trajectory-based applications.
5.12 Computation Cost
TABLE IV: Computation Cost. NA means the model does not require map-matching.
Dataset Chengdu / Porto
Methods Parameters Training Time Inference Time (seconds / batch)
(million) (seconds / epoch) Generation HMM Total
TrajVAE + HMM 12.03 217.57 / 269.56 2.79 / 2.92 128.41 / 182.58 131.20 / 185.50
TrajGAN + HMM 0.36 231.35 / 240.97 2.46 / 3.14 128.41 / 182.58 130.87 / 185.72
EETG + HMM 8.29 327.94 / 404.21 6.04 / 6.37 128.41 / 182.58 134.45 / 188.95
Diff-RNTraj 27.18 194.53 / 226.97 13.07 / 13.19 NA 13.07 / 13.19
In this section, we conduct computational cost analysis on the Chengdu and Porto datasets using a batch size of 256 with the NVIDIA A40 card. As shown in Table IV, Diff-RNTraj has a larger number of parameters compared to the baselines, as it stacks multiple RDCLs. Despite this, our model’s training time is shorter than the baselines. This efficiency is due to RDCL’s convolution-based structure, which allows for parallel processing. In contrast, TrajVAE, TrajGAN, and EETG, which are based on the LSTM architecture, process trajectories sequentially.
For the inference phase, we calculate the time cost of each model to generate a batch of trajectories. When generating trajectories, our model’s speed is slower than baselines due to the requirement of 500 iterative denoising steps. However, the baselines necessitate an additional HMM-based map-matching process after trajectory generation, which adds considerable computational expense. This process has a computational complexity of $O(T\mathcal{R}^{2})$, where $T$ represents the trajectory length and $\mathcal{R}$ is the number of road segments. Our model, on the other hand, does not require the map-matching step. Consequently, when generating road network-constrained trajectories, the total time of Diff-RNTraj still is the fastest.
6 Conclusion
@@
In real scenarios, the drive path of a vehicle usually shows different characteristics under changes in different objective factors, such as traffic conditions and departure time. In the future, we plan to study more specific trajectory generation tasks, e.g., using the different objective factors as conditions to generate trajectories under different patterns.